Allen bought 30 stamps at the post office in 39¢ and 24¢ denominations.
If the total cost of the stamps was $9.90, how many 39¢ stamps did Allen buy?

Answer by edjones(8007)

Let x=amt of .39 stamps and y=amt of .24 stamps.
A) x+y=30
B) .39x+.24y=9.9
.
x=30-y
.39(30-y)+.24y=9.9 Substitute 30-y for x in B).
11.7-.39y+.24y=9.9
-.15y+11.7=9.9
-.15y+11.7-11.7=9.9-11.7
-.15y=-1.8
-.15y/-.15=-1.8/-.15
y=12
x=30-12
x=18
Check:
.39*18+.24*12=9.9
.
Ed